The perimeter of a rectangular garden is 200 yards. The length is eight yards more than three times the width. Find the lenght of the garden.
Answer by checkley75(3666)

LENGTH=3W+8
2L+2W=200
2(3W+8)+2W=200
6W+16+2W=200
8W=200-16
8W=184
W=184/8
W=23 YARDS IS THE WIDTH.
THEREFORE THE LENGTH=3*23+8=69+8=77 YARDS FOR THE LENGTH.
PROOF
2*77+2*23=200
154+46=200
200=200